OpenStudy (anonymous):

The mass of a can is 350 grams, correct to the nearest 10 grams. Giving your answer in kilograms, calculate the lower bound for the total mass of the box and 20 identical cans

OpenStudy (anonymous):

The mass of a box is 2 kilograms, correct to the nearest kilogram

OpenStudy (rajee_sam):

the mass of the cans is correct to the nearest 10 grams meaning \[\pm 10 \space grams\]

OpenStudy (rajee_sam):

so the mass can be anywhere between 340 gm - 360 gm so 20 cans would be 20 x 340 = 6800 gm = 6.8 Kg --- Lower bound Now lower bound of mass of box + mass of 20 cans would be = 2 + 6.8 = 8.8 Kg

OpenStudy (anonymous):

The answer given is 8.4 !

OpenStudy (rajee_sam):

IDK with the given info this is what I got maybe someone else can take a look at it and see whether I did it right. @ganeshie8

OpenStudy (anonymous):

Take lower bound of each of the can which is 345. then multiply it by 20, which is 345*20 = 6900 grams. divide 6900 by 1000 you get 6.9 kg. as for the lower bound of the box it is 1.5 kg. add 6.9 and 1.5 you get 8.4